# On the Adversarial Robustness of Learning-based Conformal Novelty Detection

### v1

@@ Appendix C Proof of Theorem 1
$\text{FDR}_{\text{attack}}=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{{\widetilde{V}}_{i}}{{{\widetilde{R}}}\lor 1}\right]+\mathbb{E}\left[\sum_{i\in\mathcal{A}\cap\mathcal{H}_{0}}\frac{{{\widetilde{V}}}_{i}}{{\widetilde{R}}\lor 1}\right].$ (24)
Let ${\widetilde{R}}=\sum_{i=1}^{m}{\widetilde{V}}_{i}$ be the total number of rejections after the attack. For the second term, we can bound it by
$\mathbb{E}\left[\sum_{i\in\mathcal{A}\cap\mathcal{H}_{0}}\frac{{{\widetilde{V}}}_{i}}{{\widetilde{R}}\lor 1}\right]\leq\mathbb{E}\left[\frac{|\mathcal{A}\cap\mathcal{H}_{0}|}{{\widetilde{R}}\lor 1}\right]\overset{(a)}{=}\mathbb{E}\left[\frac{m_{\mathcal{A}}}{{\widetilde{R}}\lor 1}\right],$ (25)
where $(a)$ follows since $\mathcal{A}\subseteq\mathcal{H}_{0}$. For the first term, define $S_{i}=\tilde{s}(\widetilde{Z}_{i})$ for $i\in[1:m+n]$. Fix any $i\in\mathcal{H}_{0}\setminus\mathcal{A}$, and for $j\neq i$:
$C_{i,j}\;=\;\frac{1}{n-k+1}\left(\sum_{s\,\in\,\{S_{k+1},\dots,S_{n},\,S_{n+i}\}}\mathbf{1}\{\,s>S_{n+j}\}\right).$ (26)
@@
Hence $p^{\prime}_{j}\leq{\tilde{p}}_{j}$ whenever ${\tilde{p}}_{j}\leq{\tilde{p}}_{i}$ is guaranteed by construction ($C_{i,j}$ is smaller or equal if ${\tilde{p}}_{j}\leq{\tilde{p}}_{i}$), and $p^{\prime}_{j}={\tilde{p}}_{j}$ if ${\tilde{p}}_{j}>{\tilde{p}}_{i}$. Condition (63) in [7, Lemma D.6] is satisfied for all $j\neq i$. Recall $\tilde{\tau}$ is the BH index for $({\tilde{p}}_{1},\dots,{\tilde{p}}_{m})$ and let $\tau_{i}^{\prime}:=\tau^{\prime}_{\text{BH}}$ be the BH index for $(p^{\prime}_{1},\dots,p^{\prime}_{m})$. By [7, Lemma D.6], we obtain
@@
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{{\widetilde{V}}_{i}}{{{\widetilde{R}}}\lor 1}\right]=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{\,\mathbf{1}\{\,{\tilde{p}}_{i}\leq\alpha\,(\tilde{\tau}/m)\}}{\tilde{\tau}}\right]=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{\,\mathbf{1}\{\,{\tilde{p}}_{i}\leq\alpha\,(\tau^{\prime}_{i}/m)\}}{\tau^{\prime}_{i}}\right].$ (30)
@@
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{{\widetilde{V}}_{i}}{{{\widetilde{R}}}\lor 1}\right]$ $=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\mathbb{E}\left[\frac{\mathbf{1}\left\{{\tilde{p}}_{i}\leq\alpha\,(\tau^{\prime}_{i}/m)\right\}}{\tau^{\prime}_{i}}\,\Big|\,\widetilde{W}_{i}\right]\right]$ (32)
$=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{1}{\tau^{\prime}_{i}}\,\mathbb{E}\left[\mathbf{1}\left\{{\tilde{p}}_{i}\leq\alpha\,(\tau^{\prime}_{i}/m)\right\}\,\Big|\,\widetilde{W}_{i}\right]\right],$ (33)
@@
$\mathbb{E}\left[\mathbf{1}\{\,{\tilde{p}}_{i}\leq\alpha\,(\tau^{\prime}_{i}/m)\}\,\Big|\,\widetilde{W}_{i}\right]=\mathbb{P}\left((n-k+1)\,{\tilde{p}}_{i}\;\leq\;\alpha\,(n-k+1)\,\tfrac{\tau^{\prime}_{i}}{m}\,\Big|\,\widetilde{W}_{i}\right)=\frac{\lfloor\alpha(n-k+1)\,\tfrac{\tau^{\prime}_{i}}{m}\rfloor}{n-k+1}.$ (34)
@@
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{{\widetilde{V}}_{i}}{{{\widetilde{R}}}\lor 1}\right]=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{\lfloor\alpha(n-k+1)\,\tau^{\prime}_{i}/m\rfloor}{(n-k+1)\,\tau^{\prime}_{i}}\right]\leq m_{0}\cdot\mathbb{E}\left[\frac{\lfloor\alpha(n-k+1)\,\tau^{\prime}_{i}/m\rfloor}{(n-k+1)\,\tau^{\prime}_{i}}\right].$ (35)
@@ Appendix D Direct Decision-based Attack Scheme
  1. Data: Training samples $\{Z_{j}\}_{j=1}^{n}$ and test samples $\{Z_{j}\}_{j=n+1}^{m+n}$, but the attacker does not know which test samples are nulls and non-nulls;
  2. Algorithm: All the information about the AdaDetect implemented by the user, including the machine learning model for the score function and its parameters.
@@
  1. Step 1: Initial detection and BH labeling. Using the training data $\{Z_{j}\}_{j=1}^{n}$ and the mixed sample $\{Z_{j}\}_{j=k+1}^{n+m}$, form the dataset
@@
    Note that $s(z)$ automatically satisfies the condition in equation 3 as $Y_{i}$ for $i\in[k+1:n+m]$ are the same. Compute empirical $p$-values for $i\in[1:m]$
@@
  2. Step 2: Attack set selection. Within the set of test samples (i.e., with indices from $[n+1:n+m]$), select a subset $\{Z_{n+i}:i\in\mathcal{A}\}$ from the unrejected test samples as the attack target.
@@
  3. Step 3: Decision-based adversarial perturbation. For each $i\in\mathcal{A}$, generate
@@
  4. Step 4: Applying AdaDetect on the attacked data. After the attack, the user applies AdaDetect and computes the score function as the first step. As the data is now changed by the attacker, we denote the score function after the attack by $\tilde{s}$, and the empirical $p$-value after the attack by ${\tilde{p}}_{i}$ for $i\in[1:m]$. We stress that $\tilde{s}$ still satisfies equation 3.
@@
This proposition captures the main subtle yet important difference between this attack scheme and the oracle setting in Theorem 1. It holds because $f_{\text{attack}}(\cdot\,;\,s(z))$ only relies on the score function $s(z)$, and $s(z)$ is invariant to order of elements in $\{Z_{k+1},\dots,Z_{n+m}\}$ according to equation 3, as a consequence of the PU framework.
@@ D-A Analysis
Consider that $\mathcal{A}$ is a fixed set of indices with $m_{a}=|\mathcal{A}|$. Under Assumption 1, with the score function $\tilde{s}$ satisfying the permutation invariance property in equation 3 and the attack scheme $f_{attack}$ being order-invariant as in equation 37, the FDR after the attack is
$\text{FDR}^{*}_{\text{attack--decision}}\leq\alpha+m_{a}\cdot\mathbb{E}\left[\frac{1}{{\widetilde{R}}\lor 1}\right],$ (38)
@@
It has been proved that AdaDetect has a strong detection power (the probability of correctly rejecting a non-null), as shown in [7, Theorem 5.1]. This implies that the set $\mathcal{A}$ will nearly contain all indices from true nulls because the non-nulls are mostly rejected. We use this to show that the upper bound $\mathbb{E}[\frac{m_{\mathcal{A}}}{{\widetilde{R}}\lor 1}]$ in equation 1 is relatively tight (see details in Proposition 3).
@@
Consider that $\mathcal{A}$ is random with fixed size $|\mathcal{A}|=m_{a}$. Under Assumption 1, with equation 3 and equation 7, the FDR after the attack is
$\text{FDR}_{\text{attack}}\leq\text{FDR}^{*}_{\text{attack--decision}}\leq\alpha+m_{a}\cdot\mathbb{E}\left[\frac{1}{{\widetilde{R}}\lor 1}\right],$ (39)
where the expectations are taken over the randomness in the training and test samples $\{Z_{j}\}_{j=1}^{m+n}$, which induces randomness in $\mathcal{A}$ and $\widetilde{R}$}.
@@
Consider that $\mathcal{A}$ is random with a random size $|\mathcal{A}|=m_{\mathcal{A}}$. Under Assumption 1, with equation 3 and equation 7, the FDR after the attack is
$\text{FDR}_{\text{attack}}\leq\alpha+\mathbb{E}\left[\frac{m_{\mathcal{A}}}{{\widetilde{R}}\lor 1}\right],$ (40)
where the expectations are taken over the randomness in the training and test samples $\{Z_{j}\}_{j=1}^{m+n}$, which induces randomness in $\mathcal{A}$ and $\widetilde{R}$}.
In the following proposition, we show that the upper bound in equation 25 is relatively tight by making a connection between step (a) in equation 25 and the power of AdaDetect. Roughly speaking, the upper bound is relatively tight when the power of AdaDetect is decent. Instead of showing detailed technical steps following the proof of [7, Theorem 5.1], we choose to provide a high-level argument to connect the power of AdaDetect and our upper bound in equation 25.
@@
Under the assumptions in [7, Theorem 5.1] and when $\mathcal{A}$ is randomly selected from the unrejected indices with $m_{\mathcal{A}}=\lfloor\gamma(m-R)\rfloor$, we have that for some small $\delta^{\prime}$ and $\eta^{\prime}$,
@@
According to [7, Theorem 5.1], we have that the rejection set by AdaDetect at level $\lambda$, denoted by $\text{AdaDetect}_{\lambda}$ satisfies
@@
Since $\widetilde{\mathcal{R}}=\widetilde{\text{AdaDetect}}_{\lambda}$ and $|\widetilde{\mathcal{R}}\cap(\mathcal{H}_{1}\cup\mathcal{A})|={\widetilde{R}}-{\widetilde{V}}$, we have that the unrejected set contains $(|\mathcal{H}_{1}\cup\mathcal{A}|-({\widetilde{R}}-{\widetilde{V}})))\leq|\mathcal{H}_{1}\cup\mathcal{A}|\cdot\eta^{\prime}$ with probability at least $1-\delta^{\prime}$. Define $\mathcal{E}\;=\;\left\{\,\big|\widetilde{R}-\widetilde{V}\big|\;\geq\;|\mathcal{H}_{1}\cup\mathcal{A}|\cdot(1-\eta^{\prime})\right\}$. We use the shorthand $Z:=\{Z_{i}\}_{i=1}^{n+m}$ to denote the whole dataset. Given any fixed dataset $Z=z$, the only remaining randomness comes from the random selection (and it is independent of $Z$), while the random variables $\widetilde{R}$, $\mathcal{A}$, and $m_{\mathcal{A}}$ take on realizations as $\widetilde{R}(z)$, $\mathcal{A}(z)$, and $m_{\mathcal{A}(z)}$, respectively. We now have
@@
$\geq(1-\delta^{\prime})\;\cdot\int_{z}\frac{\displaystyle\binom{\,m-\widetilde{R}(z)-|\mathcal{H}_{1}\cup\mathcal{A}(z)|\cdot\eta^{\prime}\,}{m_{\mathcal{A}(z)}}}{\displaystyle\binom{\,m-\widetilde{R}(z)\,}{m_{\mathcal{A}(z)}}}\cdot P(Z=z|\mathcal{E})\,dz\,:=(1-\delta^{\prime})\cdot l(\eta^{\prime}),$
@@
Appendix E Additional Experiments

### v2

@@ Appendix C Proof of Theorem 1
$\hskip-1.00006pt\text{FDR}_{\text{attack}}\hskip-1.99997pt=\hskip-1.99997pt\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{{\widetilde{V}}_{i}}{{{\widetilde{R}}_{m_{a}}}\lor 1}\right]+\mathbb{E}\left[\sum_{i\in\mathcal{A}\cap\mathcal{H}_{0}}\frac{{{\widetilde{V}}}_{i}}{{\widetilde{R}}_{m_{a}}\lor 1}\right].$
Let ${\widetilde{R}}_{m_{a}}=\sum_{i=1}^{m}{\widetilde{V}}_{i}$ be the total number of rejections after the attack. For the second term, we can bound it by
$\mathbb{E}\left[\sum_{i\in\mathcal{A}\cap\mathcal{H}_{0}}\frac{{{\widetilde{V}}}_{i}}{{\widetilde{R}}_{m_{a}}\lor 1}\right]\leq\mathbb{E}\left[\frac{|\mathcal{A}\cap\mathcal{H}_{0}|}{{\widetilde{R}}_{m_{a}}\lor 1}\right]\overset{(a)}{=}\mathbb{E}\left[\frac{m_{\mathcal{A}}}{{\widetilde{R}}_{m_{a}}\lor 1}\right],$ (20)
where $(a)$ follows since $\mathcal{A}\subseteq\mathcal{H}_{0}$. For the first term, define $S_{i}=\tilde{s}(\widetilde{Z}_{i})$ for $i\in[1:m+n]$. Fix any $i\in\mathcal{H}_{0}\setminus\mathcal{A}$, and for $j\neq i$, we have
$C_{i,j}=\frac{1}{n-k+1}\left(\sum_{s\,\in\,\{S_{k+1},\dots,S_{n},\,S_{n+i}\}}\mathbf{1}\{\,s>S_{n+j}\}\right).$
@@
Hence $p^{\prime}_{j}\leq{\tilde{p}}_{j}$ whenever ${\tilde{p}}_{j}\leq{\tilde{p}}_{i}$ is guaranteed by construction ($C_{i,j}$ is smaller or equal if ${\tilde{p}}_{j}\leq{\tilde{p}}_{i}$), and $p^{\prime}_{j}={\tilde{p}}_{j}$ if ${\tilde{p}}_{j}>{\tilde{p}}_{i}$. Condition (63) in Lemma D.6 from  [7] is satisfied for all $j\neq i$. Recall $\tilde{\tau}$ is the BH index for $({\tilde{p}}_{1},\dots,{\tilde{p}}_{m})$ and let $\tau_{i}^{\prime}:=\tau^{\prime}_{\text{BH}}$ be the BH index for $(p^{\prime}_{1},\dots,p^{\prime}_{m})$. By Lemma D.6 from  [7], we obtain
@@
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{{\widetilde{V}}_{i}}{{{\widetilde{R}}_{m_{a}}}\lor 1}\right]$ $=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{\,\mathbf{1}\{\,{\tilde{p}}_{i}\leq\alpha\,(\tilde{\tau}/m)\}}{\tilde{\tau}}\right]=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{\,\mathbf{1}\{\,{\tilde{p}}_{i}\leq\alpha\,(\tau^{\prime}_{i}/m)\}}{\tau^{\prime}_{i}}\right].$
@@
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{{\widetilde{V}}_{i}}{{{\widetilde{R}}_{m_{a}}}\lor 1}\right]=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\mathbb{E}\left[\frac{\mathbf{1}\left\{{\tilde{p}}_{i}\leq\alpha\,(\tau^{\prime}_{i}/m)\right\}}{\tau^{\prime}_{i}}\,\Big|\,\widetilde{W}_{i}\right]\right]=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{1}{\tau^{\prime}_{i}}\,\mathbb{E}\left[\mathbf{1}\left\{{\tilde{p}}_{i}\leq\alpha\,(\tau^{\prime}_{i}/m)\right\}\,\Big|\,\widetilde{W}_{i}\right]\right],$
@@
$\mathbb{E}\left[\mathbf{1}\{\,{\tilde{p}}_{i}\leq\alpha\,(\tau^{\prime}_{i}/m)\}\,\Big|\,\widetilde{W}_{i}\right]=\mathbb{P}\left((n-k+1)\,{\tilde{p}}_{i}\;\leq\;\alpha\,(n-k+1)\,\tfrac{\tau^{\prime}_{i}}{m}\,\Big|\,\widetilde{W}_{i}\right)=\frac{\lfloor\alpha(n-k+1)\,\tfrac{\tau^{\prime}_{i}}{m}\rfloor}{n-k+1}.$
@@
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{{\widetilde{V}}_{i}}{{{\widetilde{R}}_{m_{a}}}\lor 1}\right]=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{\lfloor\alpha(n-k+1)\,\tau^{\prime}_{i}/m\rfloor}{(n-k+1)\,\tau^{\prime}_{i}}\right]\leq m_{0}\cdot\mathbb{E}\left[\frac{\lfloor\alpha(n-k+1)\,\tau^{\prime}_{i}/m\rfloor}{(n-k+1)\,\tau^{\prime}_{i}}\right].$
@@ Appendix D Direct Decision-based Attack Scheme
Data: Training samples $\{Z_{j}\}_{j=1}^{n}$ and test samples $\{Z_{j}\}_{j=n+1}^{m+n}$, but the attacker does not know which test samples are nulls and non-nulls;
Algorithm: All the information about the AdaDetect implemented by the user, including the machine learning model for the score function and its parameters.
@@
Step 1: Initial detection and BH labeling. Using the training data $\{Z_{j}\}_{j=1}^{n}$ and the mixed sample $\{Z_{j}\}_{j=k+1}^{n+m}$, form the dataset
@@
Note that $s(z)$ automatically satisfies the condition in equation II-A as $Y_{i}$ for $i\in[k+1:n+m]$ are the same. Compute empirical $p$-values for $i\in[1:m]$,
@@
Step 2: Attack set selection. Within the set of test samples (i.e., with indices from $[n+1:n+m]$), select a subset $\{Z_{n+i}:i\in\mathcal{A}\}$ from the unrejected test samples as the attack target.
@@
Step 3: Decision-based adversarial perturbation. For each $i\in\mathcal{A}$, generate
@@
Step 4: Applying AdaDetect on the attacked data. After the attack, the user applies AdaDetect and computes the score function as the first step. As the data is now changed by the attacker, we denote the score function after the attack by $\tilde{s}$, and the empirical $p$-value after the attack by ${\tilde{p}}_{i}$ for $i\in[1:m]$. We stress that $\tilde{s}$ still satisfies equation II-A.
@@
This proposition captures the main subtle yet important difference between this attack scheme and the oracle setting in Theorem 1. It holds because $f_{\text{attack}}(\cdot\,;\,s(z))$ only relies on the score function $s(z)$, and $s(z)$ is invariant to order of elements in $\{Z_{k+1},\dots,Z_{n+m}\}$ according to equation II-A, as a consequence of the PU framework.
@@ D-A Analysis
Consider that $\mathcal{A}$ is a fixed set of indices with $m_{a}=|\mathcal{A}|$. Under Assumption 1, with the score function $\tilde{s}$ satisfying the permutation invariance property in equation II-A and the attack scheme $f_{attack}$ being order-invariant as in equation 25, the FDR after the attack is
$\text{FDR}^{*}_{\text{attack--decision}}\leq\alpha+m_{a}\cdot\mathbb{E}\left[\frac{1}{{\widetilde{R}}_{m_{a}}\lor 1}\right],$ (26)
@@
It has been proved that AdaDetect has a strong detection power (the probability of correctly rejecting a non-null), as shown in Theorem 5.1 from [7]. This implies that the set $\mathcal{A}$ will nearly contain all indices from true nulls because the non-nulls are mostly rejected. We use this to show that the upper bound $\mathbb{E}[\frac{m_{\mathcal{A}}}{{\widetilde{R}}_{m_{a}}\lor 1}]$ in equation 1 is relatively tight (see details in Proposition 3).
@@
Consider that $\mathcal{A}$ is random with fixed size $|\mathcal{A}|=m_{a}$. Under Assumption 1, with equation II-A and equation 7, the FDR after the attack is
$\text{FDR}_{\text{attack}}\leq\text{FDR}^{*}_{\text{attack--decision}}\leq\alpha+m_{a}\cdot\mathbb{E}\left[\frac{1}{{\widetilde{R}}_{m_{a}}\lor 1}\right],$
where the expectations are taken over the randomness in the training and test samples $\{Z_{j}\}_{j=1}^{m+n}$, which induces randomness in $\mathcal{A}$ and ${\widetilde{R}}_{m_{a}}$.
@@
Consider that $\mathcal{A}$ is random with a random size $|\mathcal{A}|=m_{\mathcal{A}}$. Under Assumption 1, with equation II-A and equation 7, the FDR after the attack is
$\text{FDR}_{\text{attack}}\leq\alpha+\mathbb{E}\left[\frac{m_{\mathcal{A}}}{{\widetilde{R}}_{m_{a}}\lor 1}\right],$ (27)
where the expectations are taken over the randomness in the training and test samples $\{Z_{j}\}_{j=1}^{m+n}$, which induces randomness in $\mathcal{A}$ and ${\widetilde{R}}_{m_{a}}$.
In the following proposition, we show that the upper bound in equation 20 is relatively tight by making a connection between step (a) in equation 20 and the power of AdaDetect. Roughly speaking, the upper bound is relatively tight when the power of AdaDetect is decent. Instead of showing detailed technical steps following the proof of Theorem 5.1 from [7], we choose to provide a high-level argument to connect the power of AdaDetect and our upper bound in equation 20.
@@
Under the assumptions in Theorem 5.1 from [7] and when $\mathcal{A}$ is randomly selected from the unrejected indices with $m_{\mathcal{A}}=\lfloor\gamma(m-R)\rfloor$, we have that for some small $\delta^{\prime}$ and $\eta^{\prime}$,
@@
According to Theorem 5.1 from [7], we have that the rejection set by AdaDetect at level $\lambda$, denoted by $\text{AdaDetect}_{\lambda}$ satisfies
@@
Since $\widetilde{\mathcal{R}}=\widetilde{\text{AdaDetect}}_{\lambda}$ and $|\widetilde{\mathcal{R}}\cap(\mathcal{H}_{1}\cup\mathcal{A})|={\widetilde{R}}_{m_{a}}-{\widetilde{V}}$, we have that the unrejected set contains $(|\mathcal{H}_{1}\cup\mathcal{A}|-({\widetilde{R}}_{m_{a}}-{\widetilde{V}})))\leq|\mathcal{H}_{1}\cup\mathcal{A}|\cdot\eta^{\prime}$ with probability at least $1-\delta^{\prime}$. Define $\mathcal{E}\;=\;\left\{\,\big|\widetilde{R}-\widetilde{V}\big|\;\geq\;|\mathcal{H}_{1}\cup\mathcal{A}|\cdot(1-\eta^{\prime})\right\}$. We use the shorthand $Z:=\{Z_{i}\}_{i=1}^{n+m}$ to denote the whole dataset. Given any fixed dataset $Z=z$, the only remaining randomness comes from the random selection (and it is independent of $Z$), while the random variables $\widetilde{R}$, $\mathcal{A}$, and $m_{\mathcal{A}}$ take on realizations as $\widetilde{R}(z)$, $\mathcal{A}(z)$, and $m_{\mathcal{A}(z)}$, respectively. We now have
@@
$\hskip 5.0pt\geq(1-\delta^{\prime})\;\cdot\int_{z}\frac{\displaystyle\binom{\,m-\widetilde{R}(z)-|\mathcal{H}_{1}\cup\mathcal{A}(z)|\cdot\eta^{\prime}\,}{m_{\mathcal{A}(z)}}}{\displaystyle\binom{\,m-\widetilde{R}(z)\,}{m_{\mathcal{A}(z)}}}\cdot P(Z=z|\mathcal{E})\,dz\,:=(1-\delta^{\prime})\cdot l(\eta^{\prime}),$
@@
Appendix E Comparison between our surrogate algorithm and INCREASE-c [18]
The two approaches are fundamentally different. We focus on attacking data samples directly, while INCREASE-c perturbs p-values. From a practical perspective, the crucial difference between the two lies in the perturbation strategy. While our method targets data point near the learned decision boundary via a surrogate score function, INCREASE-c effectively seeks to alter $p$-values that are often far from the decision boundary to ensure a collective shift in the empirical FDR. Consequently, INCREASE-c requires a significantly larger per-sample perturbation to force a change in the rejection threshold, whereas our approach exploits the local sensitivity of the score function to induce misclassification with minimal data distortion. We focus on credit card dataset as in [18] and the BH threshold in our surrogate approach is about $0.0083$, and we are perturbing the data with $p$-values that are slightly above this threshold.
To evaluate the comparative impact, we replicate the INCREASE-c experimental framework on the same credit card dataset. An isolation forest is trained with a training sample selected uniformly at random from the set of true nulls; a calibration subset of strictly genuine transactions is randomly selected from the null data, after which a test sample is formed by combining the remaining null data with a random subset of fraudulent transactions. The test sample is transformed to $p$-values, and the BH procedure (with a level of $0.1$) is applied to identify the set of fraudulent transactions. (The full details of the experiment can be found in [18].) We execute INCREASE-$c$ against these $p$-values and their ground-truth labels across $100$ simulations for each attack budget $c=1,5,10,20$. The results are reported in Table $6$. The high average value of the selected $p$-values (i.e., $\mathbb{E}[p_{\text{selected}}]$) confirms that INCREASE-c targets data points far from the (average) BH cutoff (i.e., the last column in Table 6). Due to the drastic change in perturbed $p$-values, there is a clear impact on the resulting FDR and total rejection count.
TABLE VI: Experiment E.1: INCREASE-c results
$c$ original FDR INCREASE-c FDR original $\mathbb{E}[R]$ INCREASE-c $\mathbb{E}[R]$ $\mathbb{E}[p_{\text{selected}}]$ avg. BH cutoff
1 0.08 0.13 43.52 45.79 0.99 0.0052
5 0.09 0.17 55.24 64.39 0.99 0.0061
10 0.08 0.24 46.33 64.00 0.99 0.0045
20 0.08 0.31 55.48 90.28 0.98 0.0034
Appendix F Additional Experiments